OpenStudy (anonymous):

Pls help solve for x...thank you (3/cos^2x) - 7 - (4/cot x) = 0

OpenStudy (anonymous):

\[\frac{1}{\cos^2 x}=1+\tan^2 x \\ and \\ \frac{1}{\cot x}=\tan x\]

OpenStudy (lgbasallote):

\[\frac{3}{\cos^2 x} - 7 - \frac{4}{\frac{\cos x}{\sin x}} = 0\] \[\frac{3}{\cos^2 x} - 7 - \frac{4\sin x}{\cos x} = 0\] \[\frac{3}{\cos^2 x} - \frac{4\sin x}{\cos x} - 7 = 0\] \[3 \sec^2 x - 4 \tan x - 7 = 0\] \[3 (\tan^2 x + 1) - 4\tan x - 7 = 0\] \[3\tan^2 x + 3 - 4 \tan x -7 = 0\] \[3\tan^2 x - 4\tan x - 10 = 0\] let a = tan x \[3a^2 - 4a - 10 = 0\] \[(3a +5)(a - 3) = 0\] \[ a = -5/3 \; ; \; a = 3\] \[\tan x = -5/3 \;; \; \tan x = 3\] \[x = \tan^{-1} (- 5/3) \; ;\; \tan^{-1} (3)\]

OpenStudy (lgbasallote):

do you get that @kwokyy ?

OpenStudy (anonymous):

Thanks guys...